# Robust Implicit Adaptive Low Rank Time-Stepping Methods for Matrix Differential Equations

### v3

@@ Abstract
In this work, we develop implicit rank-adaptive schemes for time-dependent matrix differential equations. The dynamic low rank approximation (DLRA) is a well-known technique to capture the dynamic low rank structure based on Dirac–Frenkel time-dependent variational principle. In recent years, it has attracted a lot of attention due to its wide applicability. Our schemes are inspired by the three-step procedure used in the rank adaptive version of the unconventional robust integrator (the so called BUG integrator) [2] for DLRA. First, a prediction (basis update) step is made computing the approximate column and row spaces at the next time level. Second, a Galerkin evolution step is invoked using a base implicit solve for the small core matrix. Finally, a truncation is made according to a prescribed error threshold. Since the DLRA is evolving the differential equation projected on to the tangent space of the low rank manifold, the error estimate of the BUG integrator contains the tangent projection (modeling) error which cannot be easily controlled by mesh refinement. This can cause convergence issue for equations with cross terms.
@@ 1 Introduction
$F(X(t),t)=\sum_{j=1}^{s}A_{j}X(t)B_{j}^{T}+G(t),$
and $A_{j}\in\mathbb{R}^{m_{1}\times m_{1}},B_{j}\in\mathbb{R}^{m_{2}\times m_{2}}$ are sparse or structured matrices for which a fast matrix-vector product is assumed to be known. Further, $G(t)\in\mathbb{R}^{m_{1}\times m_{2}}$ is a given function that is assumed to have a known low rank decomposition. The integer $s,$ which denotes the separation rank of $F(\cdot),$ is assumed not to be too large. We focus on implicit methods, which are particularly needed for computing solutions to stiff equations.
@@
A rank adaptive method adaptively choose the rank at each time step according to a prescribed error tolerance. Several approaches to rank adaptivity can be found in the literature. For example, a straightforward idea is the so-called step truncation method [4] which evolves the low rank solution for one time step by a traditional time stepping method in an ambient space of higher rank, then performs a truncation (by SVD with given tolerance). This method is intuitive and effective for explicit time stepping methods. For implicit schemes [21], implementation of step truncation methods is nontrivial and requires an effective iterative solver in adaptive low rank format [7, 1]. Another approach that has been suggested by multiple groups, and that is similar to what we propose, is to modify the DLRA method by using rank increase and decrease indicators and augment or truncate the spaces at each time step [12, 11, 24]. In particular the ideas in [24], while being specialized to time dependent Schrödinger, are similar to the approach we take here. A promising idea in [2] uses the so called unconventional (BUG) integrator [3] to achieve rank adaptivity. This approach was recently extended to higher order in time in [19].
@@ 2.1 Notations and preliminaries
We use $\mathcal{T}_{\epsilon}$ to denote a generic matrix approximation operator with accuracy $\epsilon$ in the Frobenius norm, i.e. $\|A-\mathcal{T}_{\epsilon}(A)\|\leq\epsilon.$ A prominent example is the truncated SVD of a matrix. Namely, given a generic rank $r$ matrix $A\in\mathbb{R}^{m_{1}\times m_{2}}$ with reduced SVD: $A=U\Sigma V^{T},$ then $\mathcal{T}^{svd}_{\epsilon}(A)=U[:,1:s]\,{\sf diag}(\sigma_{1},\cdots,\sigma_{r})\,V[:,1:s]^{T},$ where we have used the standard MATLAB notation for submatrices and $r$ is chosen to be the smallest integer so that $(\sum_{j=s+1}^{r}\sigma_{j}^{2})^{1/2}\leq\epsilon.$
@@
Form $U=[U_{1},\ldots,U_{m}],\,\Sigma={\sf diag}(\Sigma_{1},\ldots,\Sigma_{m}),\,V=[V_{1},\ldots,V_{m}]$. Perform column pivoted QR: $[Q_{1},R_{1},\Pi_{1}]={\sf qr}(U),$ $[Q_{2},R_{2},\Pi_{2}]={\sf qr}(V).$ Compute the truncated SVD: $\mathcal{T}^{svd}_{\epsilon}(R_{1}\Pi_{1}\,\Sigma\,\Pi_{2}^{T}R_{2}^{T})=\mathcal{U}\mathcal{S}\mathcal{V}^{T}$. Form $\mathcal{U}\leftarrow Q_{1}\mathcal{U},\mathcal{V}\leftarrow Q_{2}\mathcal{V}$. Return $[\mathcal{U},\mathcal{S},\mathcal{V}]=\mathcal{T}^{sum}_{\epsilon}(\sum_{j=1}^{m}X_{j})$
@@ 2.2 Explicit step truncation schemes
(Evolution). $\hat{X}^{n+1,pre}=\hat{X}^{n}+\Delta t\mathcal{T}^{sum}_{\epsilon_{1}}(F(\hat{X}^{n},t^{n})).$ (Truncation). $\hat{X}^{n+1}=\mathcal{T}^{sum}_{\epsilon_{2}}(\hat{X}^{n+1,pre}).$
@@ 2.3 Dynamic low rank approximation
(Prediction). $K-$step and $L-$step integrating from $t^{n}$ to $t^{n+1}$. Solve
@@
to obtain $L^{n+1},$ and $[\tilde{V},\sim,\sim]={\sf qr}([V^{n},L^{n+1}]).$ (Galerkin Evolution). $S-$step: solve for $S^{n+1}$ from
@@
to obtain $S^{n+1}.$ (Truncation). $\hat{X}^{n+1}=\tilde{U}\mathcal{T}_{\epsilon_{2}}^{svd}(S^{n+1})\tilde{V}^{T}=U^{n+1}\Sigma^{n+1}(V^{n+1})^{T}.$
@@ 3.1 The Merge method
(Merge Prediction). Compute the truncated SVD of the right hand side: $[\mathcal{U},\mathcal{S},\mathcal{V}]=\mathcal{T}^{sum}_{\epsilon_{1}}(F(\hat{X}^{n},t^{n})),$. Compute the BUG prediction spaces $K^{n+1},L^{n+1}$ according to Algorithm 3. Orthogonalize the merged spaces by column pivoted QR to get the prediction spaces $[\tilde{U},\sim,\sim]={\sf qr}([[\hat{U}^{n},\mathcal{U},K^{n+1}]),$ $[\tilde{V},\sim,\sim]={\sf qr}([\hat{V}^{n},\mathcal{V}^{n},L^{n+1}]).$ (Galerkin Evolution). Find $\tilde{\Sigma}$ by solving $\tilde{\Sigma}=\tilde{U}^{T}(\hat{X}^{n}+\Delta tF(\tilde{U}\tilde{\Sigma}\tilde{V}^{T},t^{n+1}))\tilde{V}.$ (Truncation). $\hat{X}^{n+1}=\hat{U}^{n+1}\hat{\Sigma}^{n+1}(\hat{V}^{n+1})^{T}=\tilde{U}\mathcal{T}_{\epsilon_{2}}^{svd}(\tilde{\Sigma})\tilde{V}^{T}.$
@@
    In the iteration we use a dense Sylvester solve with cost $\mathcal{O}((rs)^{3})$ for each iteration. When the matrices $A_{1}$ and $B_{2}$ correspond to approximations to second derivatives in the 1 and 2 direction and $A_{2}$ and $B_{1}$ are identity or diagonal positive definite matrices we find that this iteration converges to machine precision in a handful of iterations.
@@ 3.2 The Merge-adapt method
As will be shown in Lemma 3.3, the column and row spaces generated from the current solution $\hat{U}^{n},\hat{V}^{n}$ combined with $\mathcal{U}$ and $\mathcal{V}$ in $[\mathcal{U},\mathcal{S},\mathcal{V}]=\mathcal{T}^{sum}_{\epsilon_{1}}(F(\hat{X}^{n},t^{n})),$ will be sufficiently accurate to yield a first order accurate solution as long as the time step is sufficiently small compared to the Lipschitz constant $L$ of $\|F\|$. This time step restriction can be overcome by merging with the BUG spaces $K^{n+1},L^{n+1}$, which are computed useing an implicit solver and can handle the stiffness as shown in the Merge method in Algorithm 4. However, since the computation of $K^{n+1}$ and $L^{n+1}$ requires a linear solve, it is preferable to only add those when necessary. This motivates the design of the Merge-adapt method, which is described in Algorithm 5. In Steps 1-4 in Algorithm 5, we perform a calculation based purely on column and row spaces generated by the explicit scheme for the term $\mathcal{T}^{sum}_{\epsilon_{1}}(F(\hat{X}^{n},t^{n})).$ In Step 5, we do a residual check, and if this fails, the method will fall back to the Merge method as shown in Steps 6-8. Otherwise, we proceed to the next time step.
@@
The low rank solution is not, in general, expected to solve this equation but we may introduce the residual $\hat{R}^{n+1}$ so that
@@
(Cheap Prediction). Compute a first order prediction of the column and row spaces of $\hat{X}^{n+1}$. Compute the truncated SVD of the right hand side: $[\mathcal{U},\mathcal{S},\mathcal{V}]=\mathcal{T}^{sum}_{\epsilon_{1}}(F(\hat{X}^{n},t^{n}))$. Orthogonalize by column pivoted QR to get the prediction spaces $[\tilde{U},\sim,\sim]={\sf qr}([[\hat{U}^{n},\mathcal{U}]),$ $[\tilde{V},\sim,\sim]={\sf qr}([\hat{V}^{n},\mathcal{V}^{n}]).$ (Galerkin Evolution). Find $\tilde{\Sigma}$ by solving $\tilde{\Sigma}=\tilde{U}^{T}(\hat{X}^{n}+\Delta tF(\tilde{U}\tilde{\Sigma}\tilde{V}^{T},t^{n+1}))\tilde{V}.$ (Truncation). $\hat{X}^{n+1}=\hat{U}^{n+1}\hat{\Sigma}^{n+1}(\hat{V}^{n+1})^{T}=\tilde{U}\mathcal{T}_{\epsilon_{2}}^{svd}(\tilde{\Sigma})\tilde{V}^{T}.$ (Residual Check). Compute $\hat{R}^{n+1}=\hat{X}^{n+1}-\hat{X}^{n}-\Delta tF(\hat{X}^{n+1},t^{n+1})$. If $\|\hat{R}^{n+1}\|<\epsilon_{2}$ return the solution $\hat{X}^{n+1}$. If not (Merge Predicion). Compute the BUG prediction spaces $K^{n+1},L^{n+1}$ according to Algorithm 3. Orthogonalize the merged spaces by column pivoted QR: $[\tilde{U},\sim,\sim]={\sf qr}([[\hat{U}^{n},\mathcal{U},K^{n+1}]),$ $[\tilde{V},\sim,\sim]={\sf qr}([\hat{V}^{n},\mathcal{V}^{n},L^{n+1}]).$ (Galerkin Evolution). Find $\tilde{\Sigma}$ by solving $\tilde{\Sigma}=\tilde{U}^{T}(\hat{X}^{n}+\Delta tF(\tilde{U}\tilde{\Sigma}\tilde{V}^{T},t^{n+1}))\tilde{V}.$ (Truncation). $\hat{X}^{n+1}=\hat{U}^{n+1}\hat{\Sigma}^{n+1}(\hat{V}^{n+1})^{T}=\tilde{U}\mathcal{T}_{\epsilon_{2}}^{svd}(\tilde{\Sigma})\tilde{V}^{T}.$
@@ 3.3.1 Stability
[…]
@@ 4 Numerical results
We now present numerical results that illustrate the features of the methods described above. In this section we will refer to the different methods in figures and tables according to the following naming convention. We will denote the Merge method by M, and the Merge-adapt method by MA. For many problems we will also compare to the classic implicit Euler discretization as IE or Implicit Euler. When computing errors we will compare to a solution computed using Matlab’s built in ODE15s or ODE45 solvers with relative and absolute tolerance set to $10^{-12}$. We always report relative errors in the Frobenius norm. We note that in all the examples below we use $\epsilon_{1}=0$ but that the results are very similar for $\epsilon_{1}=\Delta t$.
@@
$\frac{\partial\rho}{\partial t}+r_{1}(x_{1})\frac{\partial\rho}{\partial x_{2}}+r_{2}(x_{2})\frac{\partial\rho}{\partial x_{1}}=\\ b_{1}(x_{2})\frac{\partial}{\partial x_{1}}\left[a_{1}(x_{1})\frac{\partial\rho}{\partial x_{1}}\right]+b_{2}(x_{2})\frac{\partial^{2}\left[a_{2}(x_{1})\rho\right]}{\partial x_{1}\partial x_{2}}+a_{3}(x_{1})\frac{\partial^{2}\left[b_{3}(x_{2})\rho\right]}{\partial x_{1}\partial x_{2}}+a_{4}(x)\frac{\partial}{\partial y}\left[b_{4}(y)\frac{\partial\rho}{\partial y}\right].$ (12)
@@
$\frac{\partial}{\partial x_{1}}\left[a_{2}(x_{1})\frac{\partial\rho(t_{n},x_{1},x_{1})}{\partial x_{2}}\right]\Bigg{|}_{x_{1,i},x_{2,j}}\approx D_{0}^{1}a_{1}(x_{1,i})D_{0}^{2}\,X_{i,j}^{n}.$

### v4

@@ Abstract
In this work, we develop implicit rank-adaptive schemes for time-dependent matrix differential equations. The dynamic low rank approximation (DLRA) is a well-known technique to capture the dynamic low rank structure based on Dirac–Frenkel time-dependent variational principle. In recent years, it has attracted a lot of attention due to its wide applicability. Our schemes are inspired by the three-step procedure used in the rank adaptive version of the unconventional robust integrator (the so called BUG integrator) [2] for DLRA. First, a prediction (basis update) step is made computing the approximate column and row spaces at the next time level. Second, a Galerkin evolution step is invoked using an implicit solves for the small core matrix. Finally, a truncation is made according to a prescribed error threshold. Since the DLRA is evolving the differential equation projected on to the tangent space of the low rank manifold, the error estimate of the BUG integrator contains the tangent projection (modeling) error which cannot be easily controlled by mesh refinement. This can cause convergence issue for equations with cross terms.
@@ 1 Introduction
$F(X(t),t)=\sum_{j=1}^{s_{\rm sep}}A_{j}X(t)B_{j}^{T}+G(t),$
and $A_{j}\in\mathbb{R}^{m_{1}\times m_{1}},B_{j}\in\mathbb{R}^{m_{2}\times m_{2}}$ are sparse or structured matrices for which a fast matrix-vector product is assumed to be known. Further, $G(t)\in\mathbb{R}^{m_{1}\times m_{2}}$ is a given function that is assumed to have a known low rank decomposition. The integer ${s_{\rm sep}}$, which denotes the separation rank of $F(\cdot),$ is assumed not to be too large. We focus on implicit methods, which are particularly needed for computing solutions to stiff equations.
@@
A rank adaptive method adaptively choose the rank at each time step according to a prescribed error tolerance. Several approaches to rank adaptivity can be found in the literature. For example, a straightforward idea is the so-called step truncation method [4] which evolves the low rank solution for one time step by a traditional time stepping method in an ambient space of higher rank, then performs a truncation (by SVD with given tolerance). This method is intuitive and effective for explicit time stepping methods. For implicit schemes [22], implementation of step truncation methods is nontrivial and requires an effective iterative solver in adaptive low rank format [7, 1]. Another approach that has been suggested by multiple groups, and that is similar to what we propose, is to modify the DLRA method by using rank increase and decrease indicators and augment or truncate the spaces at each time step [12, 11, 26]. In particular the ideas in [26], while being specialized to time dependent Schrödinger, are similar to the approach we take here. A promising idea in [2] uses the so called unconventional (BUG) integrator [3] to achieve rank adaptivity. This approach was recently extended to higher order in time in [20]. Finally we note the recent paper, [13], that uses the extended Krylov subspace method, [24] to construct adaptive-rank implicit time integrators.
@@ 2.1 Notations and preliminaries
We use $\mathcal{T}_{\epsilon}$ to denote a generic matrix approximation operator with accuracy $\epsilon$ in the Frobenius norm, i.e. $\|A-\mathcal{T}_{\epsilon}(A)\|\leq\epsilon.$ A prominent example is the truncated SVD of a matrix. Namely, given a generic rank $r$ matrix $A\in\mathbb{R}^{m_{1}\times m_{2}}$ with reduced SVD: $A=U\Sigma V^{T},$ then $\mathcal{T}^{svd}_{\epsilon}(A)=U[:,1:s]\,{\sf diag}(\sigma_{1},\cdots,\sigma_{s})\,V[:,1:s]^{T},$ where we have used the standard MATLAB notation for submatrices and ${r}$ is chosen to be the smallest integer so that $(\sum_{j=s+1}^{r}\sigma_{j}^{2})^{1/2}\leq\epsilon.$
@@
2 Form $U=[U_{1},\ldots,U_{m}],\,\Sigma={\sf diag}(\Sigma_{1},\ldots,\Sigma_{m}),\,V=[V_{1},\ldots,V_{m}]$.
3Perform column pivoted QR: $[Q_{1},R_{1},\Pi_{1}]={\sf qr}(U),$ $[Q_{2},R_{2},\Pi_{2}]={\sf qr}(V).$
4Compute the truncated SVD: $\mathcal{T}^{svd}_{\epsilon}(R_{1}\Pi_{1}\,\Sigma\,\Pi_{2}^{T}R_{2}^{T})=\mathcal{U}\mathcal{S}\mathcal{V}^{T}$.
5Form $\mathcal{U}\leftarrow Q_{1}\mathcal{U},\mathcal{V}\leftarrow Q_{2}\mathcal{V}$.
Return $[\mathcal{U},\mathcal{S},\mathcal{V}]=\mathcal{T}^{sum}_{\epsilon}(\sum_{j=1}^{m}X_{j})$
@@ 2.2 Explicit step truncation schemes
2 (Evolution). $\hat{X}^{n+1,pre}=\hat{X}^{n}+\Delta t\mathcal{T}^{sum}_{\epsilon_{1}}(F(\hat{X}^{n},t^{n})).$
(Truncation). $\hat{X}^{n+1}=\mathcal{T}^{sum}_{\epsilon_{2}}(\hat{X}^{n+1,pre}).$
@@ 2.3 Dynamic low rank approximation
1 (Prediction). $K-$step and $L-$step integrating from $t^{n}$ to $t^{n+1}$. Solve
@@
to obtain $L^{n+1},$ and $[\tilde{V},\sim,\sim]={\sf qr}([V^{n},L^{n+1}]).$
2(Galerkin Evolution). $S-$step: solve for $S^{n+1}$ from
@@
to obtain $S^{n+1}.$
3(Truncation). $\hat{X}^{n+1}=\tilde{U}\mathcal{T}_{\epsilon_{2}}^{svd}(S^{n+1})\tilde{V}^{T}=U^{n+1}\Sigma^{n+1}(V^{n+1})^{T}.$
@@ 3.1 The Merge method
2 (Merge Prediction). Compute the truncated SVD of the right hand side: $[\mathcal{U},\mathcal{S},\mathcal{V}]=\mathcal{T}^{sum}_{\epsilon_{1}}(F(\hat{X}^{n},t^{n})),$.
3Compute the BUG prediction spaces $K^{n+1},L^{n+1}$ according to Algorithm 3.
4Orthogonalize the merged spaces by column pivoted QR to get the prediction spaces $[\tilde{U},\sim,\sim]={\sf qr}([[\hat{U}^{n},\mathcal{U},K^{n+1}]),$ $[\tilde{V},\sim,\sim]={\sf qr}([\hat{V}^{n},\mathcal{V}^{n},L^{n+1}]).$
5(Galerkin Evolution). Find $\tilde{\Sigma}$ by solving $\tilde{\Sigma}=\tilde{U}^{T}(\hat{X}^{n}+\Delta tF(\tilde{U}\tilde{\Sigma}\tilde{V}^{T},t^{n+1}))\tilde{V}.$
(Truncation). $\hat{X}^{n+1}=\hat{U}^{n+1}\hat{\Sigma}^{n+1}(\hat{V}^{n+1})^{T}=\tilde{U}\mathcal{T}_{\epsilon_{2}}^{svd}(\tilde{\Sigma})\tilde{V}^{T}.$
@@
    In the iteration we use a dense Sylvester solver with cost $\mathcal{O}((rs)^{3})$ for each iteration. When the matrices $A_{1}$ and $B_{2}$ correspond to approximations to second derivatives in the 1 and 2 direction and $A_{2}$ and $B_{1}$ are identity or diagonal positive definite matrices we find that this iteration converges to machine precision in a handful of iterations.
@@ 3.2 The Merge-adapt method
As will be shown in Lemma 3.3, the column and row spaces generated from the current solution $\hat{U}^{n},\hat{V}^{n}$ combined with $\mathcal{U}$ and $\mathcal{V}$ in $[\mathcal{U},\mathcal{S},\mathcal{V}]=\mathcal{T}^{sum}_{\epsilon_{1}}(F(\hat{X}^{n},t^{n})),$ will be sufficiently accurate to yield a first order accurate solution as long as the time step is sufficiently small compared to the Lipschitz constant $L$ of $\|F\|$. This time step restriction can be overcome by merging with the BUG spaces $K^{n+1},L^{n+1}$, which are computed useing an implicit solver and can handle the stiffness as shown in the Merge method in Algorithm 4. However, since the computation of $K^{n+1}$ and $L^{n+1}$ requires solving a linear system, it is preferable to only add those when necessary. This motivates the design of the Merge-adapt method, which is described in Algorithm 5. In Steps 1-4 in Algorithm 5, we perform a calculation based purely on column and row spaces generated by the explicit scheme for the term $\mathcal{T}^{sum}_{\epsilon_{1}}(F(\hat{X}^{n},t^{n})).$ In Step 5, we do a residual check, and if this fails, the method will fall back to the Merge method as shown in Steps 6-8. Otherwise, we proceed to the next time step.
@@
Here $\otimes$ denotes Kronecker product and we have introduced the matrix $C$ for notational convenience The low rank solution is not, in general, expected to solve this equation but we may introduce the residual $\hat{R}^{n+1}$ so that
@@
1 (Cheap Prediction). Compute a first order prediction of the column and row spaces of $\hat{X}^{n+1}$. Compute the truncated SVD of the right hand side: $[\mathcal{U},\mathcal{S},\mathcal{V}]=\mathcal{T}^{sum}_{\epsilon_{1}}(F(\hat{X}^{n},t^{n}))$.
2Orthogonalize by column pivoted QR to get the prediction spaces $[\tilde{U},\sim,\sim]={\sf qr}([[\hat{U}^{n},\mathcal{U}]),$ $[\tilde{V},\sim,\sim]={\sf qr}([\hat{V}^{n},\mathcal{V}^{n}]).$
3(Galerkin Evolution). Find $\tilde{\Sigma}$ by solving $\tilde{\Sigma}=\tilde{U}^{T}(\hat{X}^{n}+\Delta tF(\tilde{U}\tilde{\Sigma}\tilde{V}^{T},t^{n+1}))\tilde{V}.$
4(Truncation). $\hat{X}^{n+1}=\hat{U}^{n+1}\hat{\Sigma}^{n+1}(\hat{V}^{n+1})^{T}=\tilde{U}\mathcal{T}_{\epsilon_{2}}^{svd}(\tilde{\Sigma})\tilde{V}^{T}.$
5(Residual Check). Compute $\hat{R}^{n+1}=\hat{X}^{n+1}-\hat{X}^{n}-\Delta tF(\hat{X}^{n+1},t^{n+1})$. If $\|\hat{R}^{n+1}\|<\epsilon_{2}$ return the solution $\hat{X}^{n+1}$. If not
6(Merge Predicion). Compute the BUG prediction spaces $K^{n+1},L^{n+1}$ according to Algorithm 3. Orthogonalize the merged spaces by column pivoted QR: $[\tilde{U},\sim,\sim]={\sf qr}([[\hat{U}^{n},\mathcal{U},K^{n+1}]),$ $[\tilde{V},\sim,\sim]={\sf qr}([\hat{V}^{n},\mathcal{V}^{n},L^{n+1}]).$
7(Galerkin Evolution). Find $\tilde{\Sigma}$ by solving $\tilde{\Sigma}=\tilde{U}^{T}(\hat{X}^{n}+\Delta tF(\tilde{U}\tilde{\Sigma}\tilde{V}^{T},t^{n+1}))\tilde{V}.$
8(Truncation). $\hat{X}^{n+1}=\hat{U}^{n+1}\hat{\Sigma}^{n+1}(\hat{V}^{n+1})^{T}=\tilde{U}\mathcal{T}_{\epsilon_{2}}^{svd}(\tilde{\Sigma})\tilde{V}^{T}.$
@@ 3.3.1 Stability
$\|\hat{X}^{n+1}\|\leq e^{\alpha\Delta t}\|\hat{X}^{n}\|,\quad\|\hat{X}^{n}\|\leq e^{\alpha\,t^{n}}\|\hat{X}^{0}\|,\quad\textrm{if}\,\,\alpha\Delta t\leq 1.$
@@ 4 Numerical results
We now present numerical results that illustrate the features of the methods described above. In this section we will refer to the different methods in figures and tables according to the following naming convention. We will denote the Merge method by M, and the Merge-adapt method by MA. For many problems we will also compare to the classic implicit Euler discretization as IE or Implicit Euler. When computing errors we will compare to a solution computed using Matlab’s built in ODE15s or ODE45 solvers with relative and absolute tolerance set to $10^{-12}$. We always report relative errors in the Frobenius norm. We note that in all the examples below we use $\epsilon_{1}=0$ but that the results are very similar for $\epsilon_{1}=\Delta t$. Additionally, we adjust the truncation $\epsilon_{2}$ to the PDE setting and chose $\epsilon_{2}=\mathcal{O}((\Delta t^{2}+h_{1}^{3}+h_{2}^{3})/\sqrt{h_{1}h_{2}}).$ Here the first term is from the time stepping local truncation error, the second term is from the spatial discretization error and the denominator is rescaling for the norms to be independent of grid spacing. In some of the results we also include the “hyperbolic and parabolic CFL numbers”, ($\Delta t/h$ and $\Delta t/h^{2}$) to illustrate that the method can take large timesteps when parabolic terms are present while still achieving temporal accuracy.
@@
$\frac{\partial\rho}{\partial t}+r_{1}(x_{1})\frac{\partial\rho}{\partial x_{2}}+r_{2}(x_{2})\frac{\partial\rho}{\partial x_{1}}=\\ b_{1}(x_{2})\frac{\partial}{\partial x_{1}}\left[a_{1}(x_{1})\frac{\partial\rho}{\partial x_{1}}\right]+b_{2}(x_{2})\frac{\partial^{2}\left[a_{2}(x_{1})\rho\right]}{\partial x_{1}\partial x_{2}}+a_{3}(x_{1})\frac{\partial^{2}\left[b_{3}(x_{2})\rho\right]}{\partial x_{1}\partial x_{2}}+a_{4}(x_{1})\frac{\partial}{\partial x_{2}}\left[b_{4}(x_{2})\frac{\partial\rho}{\partial x_{2}}\right].$ (12)
@@
$\frac{\partial^{2}\left[a_{2}(x_{1})\rho\right]}{\partial x_{1}\partial x_{2}}\Bigg{|}_{x_{1,i},x_{2,j}}=\frac{\partial}{\partial x_{1}}\left[a_{2}(x_{1})\frac{\partial\rho(t_{n},x_{1},x_{2})}{\partial x_{2}}\right]\Bigg{|}_{x_{1,i},x_{2,j}}\approx D_{0}^{1}a_{2}(x_{1,i})D_{0}^{2}\,X_{i,j}^{n}.$
@@
The equation (12) thus take the semidiscrete form
$\frac{d}{dt}X+R_{1}X(D_{0}^{(2)})^{T}+D_{0}^{(1)}X(R_{2})^{T}=\\ L_{a_{1}}XB_{1}^{T}+D_{0}^{(1)}A_{2}X(D_{0}^{(2)})^{T}B_{2}^{T}+A_{3}D_{0}^{(1)}XB_{3}^{T}(D_{0}^{(2)})^{T}+A_{4}XL_{b_{4}}^{T}.$ (13)
Here $R_{1},R_{2},B_{1},A_{2},B_{2},A_{3},B_{4},A_{4}$ are diagonal matrices with $r_{1}(x_{1}),r_{2}(x_{2}),b_{1}(x_{2}),$ etc. on the diagonal. Their sizes are $n_{1}\times n_{1}$ or $n_{2}\times n_{2}$ depending on if they appear on the left or right of $X$. The matrices $D_{0}^{(k)},\,k=1,2,$ are matrices with super-diagonal $0.5/h_{k}$ and sub-diagonal $-0.5/h_{k}$. The matrices $L_{a_{1}}$ and $L_{b_{4}}$ have elements $(L_{a_{1}})_{i,i-1}=(a_{1}(x_{1,i})+a_{1}(x_{1,i-1}))/(2h_{1}^{2})$, $(L_{a_{1}})_{i,i-1}=(a_{1}(x_{1,i+1})+2a_{1}(x_{1,i})+a_{1}(x_{1,i-1}))/(4h_{1}^{2})$, and $(L_{a_{1}})_{i,i+1}=(a_{1}(x_{1,i+1})+a_{1}(x_{1,i}))/(2h_{1}^{2}),$ (with $L_{b_{4}}$) defined similarly.